I have a hexagon and its side lengths are of 14cm, and it has a circular bulls eye right in the muddle with a diameter of 3cm. what is the probability that the dart that hits the target will hit the bulls eye?

Answer by Boreal(15235)

The area of a hexagon is 3 sqrt(3)/2)*a^2, where a is the side length, but if one doesn't know that off hand) there are 12 30-60-90 triangles with side length 7 and height 7 sqrt*(3). There are therefore 6 triangles with side length 14 and height 7 sqrt(3).
The area of each is 49 sqrt(3), and the whole thing therefore is 294 sqrt(3) cm ^2.
The whole in the center has area π*r^2 cm^2=(2.25*π cm^2)=7.07 cm^2.
The probability it will hit that circle is the ratio of 7.07 to 294 sqrt(3)=0.0139